OpenStudy (anonymous):

A function is shown below: f(x) = x3 + 3x2 - x - 3 What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x).

OpenStudy (anonymous):

Use Demos Graphing Calculator it is the best help

OpenStudy (triciaal):

one approach f(0) = -3 constant = 3. what are the factors of -3: 1,3, -1, -3 find f(x) using these values when f(x) = 0 you have the intercepts f(1) =(1)^3 + 3(1)^2 -(1) -3 = 1 + 3 -1 -3 = 0 (x-1) is a factor of x3 + 3x2 - x - 3 when you divide and factor points are (1,0), (0, -3), (-3.0) (-1,0)

OpenStudy (triciaal):

|dw:1408147423168:dw

OpenStudy (triciaal):